OpenStudy (anonymous):

State the slope and y-intercept of each line with the given equation and graph 13.) 2x-3y=7 14.)x=-3

OpenStudy (anonymous):

13.) 2x-3y=7 slope= 2/3 y-intercept y= 7/3 14.)x=-3 slope undefined y- intercept : none

OpenStudy (anonymous):

13). 2x-3y=7 => -3y=7-2x => y=(7/-3) +(2/3)x =>y intercept therefore, slope =2/3 14),x=-3 =>(1/-3)x =0 =>y=0=>y intercept =>(-1/3)x =y slope= (-1/3)